Angela has $90 to spend on movies and books. The price of movies (M) is $10 and the price of books (B) is $15. Which of the following one or more equations describe Angela's budget constraint?

A. 9/M = 1+ (3B/2M)

B. 10M + 15B = 90

C. B = 6 + (2/3)M

D. 2M = 18 - 3B10. Angela has $90 to spend on movies and books. The price of movies (M) is $10 and the price of books (B) is $15. Which of the following one or more equations describe Angelas budget constraint? 9/M 1+ (3B/2M) 10M 15B90 В 6 + (2/3JM 2M- 18-3B
Answer #1

Budget constraints om t C B = 90 Divide the equation by a 2 m +3B = 18 2m = 18-3B am+3B = 18 3. B = 18-2m B - 6 - 2 m 3 9 = m